# arXiv:2603.00901v1[astro-ph.SR] 01 Mar 2026

### v1

@@ 1 Introduction
Using WDLF determinations based on white-dwarf samples from the Sloan Digital Sky Survey (SDSS) and the SuperCOSMOS Sky Survey (SSS), [51] reported an upper bound $g_{ae}<2.3\times 10^{-13}$ at 95% confidence level. Moreover, analyses of these earlier WDLFs [51, 42] found a modest improvement in the fit for axion masses in the range $2.5<m_{a}\cos^{2}\beta<7.5~\mathrm{meV}$. It is also worth noting that the WDLF employed in these analyses are based on magnitude-limited samples, using the $1/V_{max}$ method [71] to account for the incompleteness of the sample at low luminosities [35, 24]. In this method, each star is counted weighted by $1/V_{max}$, where $V_{max}$ is the maximum volume where the star would still be visible considering the magnitude cut of the sample. This way, fainter stars contribute more to the WDLF than brighter stars. Although this method has been widely used in past decades, recent advances in the observed population of white dwarfs have allowed the construction of volume-limited samples. Indeed, the Gaia space mission has provided a 100 pc volume-limited sample that reaches $\sim 94\%$ for a parallax relative error lower than $\sim 10\%$ [46] . Once a proper classification of the distinct Galactic populations [76] and spectral classifications are performed [45, 33, 84] this sample allows the construction of a WDLF exempt from typical observational biases present in magnitude-limited samples [see e.g. 45, 82]. Furthermore, these previous analyses of axion properties using WDLFs adopted proper-motion selection methods, while Gaia parallax measurements have enabled the determination of distances and absolute magnitudes for these stars. Also, new classification techniques based on Random Forest algorithms that use the full parameter space of the observations are now available to separate those white dwarfs that belong to the Galactic thin disk [76].
@@ 2.3 Initial models and white dwarf model grid
Figure 1: Evolutionary tracks of white dwarfs with $Z=0.01$ and masses in the range $0.48–1.05\;M_{\odot}$. Each panel considers different values of $g_{ae}$. The color scale shows the intensity of axion emission relative to the surface luminosity.
@@ 3.1 Monte Carlo population synthesis code
To construct the theoretical WDLFs for the Galactic disk we made use of the Monte Carlo population synthesis code [32] for synthetic white dwarf populations. This code has been widely used in the analysis of single [e.g. 79, 80, 46, 81] and binary [e.g. 9, 18, 10, 75] populations of white dwarfs, as in studies of open and globular clusters [e.g 31, 78] and the Galactic bulge [77]. A detailed description of the code can be found in these references. Therefore, we will only review its basic inputs.
@@ 4 Impact of the axion emission in the WDLF
Figure 6: WDLFs in the $9<M_{G}<11.5$ range for different values of $g_{ae}$. In black and red dots the observed values, using Garcia-Zamora et al. DA sample [33] and —normalized to [33]— the Jimenez-Esteban et al. [45] DA sample, respectively.
@@ 5 Conclusions and future work
In fact, our results are nearly a factor of two less constraining than the recent determination of [30], who analyzed white-dwarf cooling in the globular cluster 47 Tuc. This difference is consistent with our expectations: the dominant uncertainty in the Galactic-disk WDLF arises from the star-formation history, it seems natural that a stellar population with well-constrained distance, star-formation rate, and metallicity would yields stronger limits. Nevertheless, 47 Tuc represents a single cluster, and the resulting inference may be sensitive to cluster-specific systematics. In particular, the observed white-dwarf sequence is derived from specific regions of the outer field of 47 Tuc [47] and may be affected by position-dependent effects such as crowding and mass segregation. More generally, the unavoidable presence of unknown unknowns is a key consideration when assessing the robustness of astrophysical constraints on particle physics. These issues highlight the complementarity between field—and cluster—based approaches. Likewise, our bounds are comparable to, but complementary with, those derived from the TRGB [83, 11], which are subject to a very distinct set of systematic uncertainties. Besides the stellar formation rate, other limitations of this work include the neglect of mergers and binary evolution, and—on a more minor note—the adopted Galactic scale height, and the omission of axion effects on stellar thermal structure from the main sequence onward. In this context, we note that [28] report that ALP couplings can affect the estimated ages of massive CO white dwarf progenitors and modify the IFMR, potentially influencing population synthesis studies. However, these effects are not relevant for our purposes, as the disk ages explored here are substantially larger than the variations introduced by these ALP-induced corrections.
@@ Appendix A Effects of stellar formation bursts on the WDLF
For instance, in Fig. 8 we show the WDLFs in the range $9<M_{G}<13$ for different values of $g_{ae}$, assuming a constant SFR, together with the corresponding $p$-values as a function of $g_{ae}$. Although the fits systematically worsen for larger values of the axion–electron coupling $g_{ae}$, none of the tests yields a good fit. Normalizing the counts over this magnitude range leads to an underestimation of the WDLF at $M_{G}\lesssim 11.5$, followed by an overestimation at the fainter end.
@@
Indeed, we find that it is possible to reproduce the observed WDLF morphology by adopting a SFR with additional degrees of freedom. In particular, we obtain acceptable fits ($p$-value $\gtrsim 74$) for a model consisting of a constant SFR over 10.6 Gyr plus two Gaussian bursts (see Fig. 9). However, this model introduces several free parameters, including the mean age, width, and normalization of each burst, as well as the normalization of the constant component. This highlights a fundamental limitation of this approach: In order to use the WDLF to robustly constrain the axion–electron coupling, the SFR must be determined independently of the white dwarf sample; otherwise, the additional free parameters could be used to artificially improve or degrade the quality of the fits for different values of $g_{ae}$. Given the current uncertainties in the observational determination of the SFR in the solar neighborhood, this requirement appears difficult to satisfy.
@@
Figure 9: Left panel: WDLF adjustment using three constant SFR components. A main component of $10.6$ Gyr and two bell-shaped bursts, resulting in a $p$-value $\gtrsim 74\%$. Right panel: the adjusted SFR.
@@ Acknowledgments
PIP-2971 and PICT 2020-03316. MEC acknowledges grant RYC2021-032721-I, funded by MCIN/AEI/10.13039
/501100011033.   MEC  and  ST  were  supported  by  the   European  UnionNextGeneration
EU/PRTR, the AGAUR/Generalitat de Catalunya grant SGR-386/2021, and by the Spanish MINECO grant PID2023-148661NB-I00. AC is supported by an ERC STG grant (“AstroDarkLS”, grant No. 101117510). AC acknowledges the Weizmann Institute of Science for hospitality at different stages of this project and the support from the Benoziyo Endowment Fund for the Advancement of Science. MEC and ST acknowledge Jordi Isern for fruitful discussions.

### v2

# New axion bounds derived from the 100-parsec Gaia DR3 white dwarf luminosity function

@@ 1 Introduction
Using WDLF determinations based on white-dwarf samples from the Sloan Digital Sky Survey (SDSS) and the SuperCOSMOS Sky Survey (SSS), [51] reported an upper bound $g_{ae}<2.3\times 10^{-13}$ at 95% confidence level. Moreover, analyses of these earlier WDLFs [51, 42] found a modest improvement in the fit for axion masses in the range $2.5<m_{a}\cos^{2}\beta<7.5~\mathrm{meV}$. It is also worth noting that the WDLF employed in these analyses are based on magnitude-limited samples, using the $1/V_{max}$ method [72] to account for the incompleteness of the sample at low luminosities [35, 24]. In this method, each star is counted weighted by $1/V_{max}$, where $V_{max}$ is the maximum volume where the star would still be visible considering the magnitude cut of the sample. This way, fainter stars contribute more to the WDLF than brighter stars. Although this method has been widely used in past decades, recent advances in the observed population of white dwarfs have allowed the construction of volume-limited samples. Indeed, the Gaia space mission has provided a 100 pc volume-limited sample that reaches $\sim 94\%$ for a parallax relative error lower than $\sim 10\%$ [46] . Once a proper classification of the distinct Galactic populations [77] and spectral classifications are performed [45, 33, 85] this sample allows the construction of a WDLF exempt from typical observational biases present in magnitude-limited samples [see e.g. 45, 83]. Furthermore, these previous analyses of axion properties using WDLFs adopted proper-motion selection methods, while Gaia parallax measurements have enabled the determination of distances and absolute magnitudes for these stars. Also, new classification techniques based on Random Forest algorithms that use the full parameter space of the observations are now available to separate those white dwarfs that belong to the Galactic thin disk [77].
@@ 2.3 Initial models and white dwarf model grid
Figure 1: Evolutionary tracks in the color-magnitude diagram of white dwarfs with metallicity $Z=0.01$, and masses in the range $0.48–1.05\;M_{\odot}$. Each panel considers different values of $g_{ae}$. The color scale shows the intensity of axion emission relative to the surface luminosity.
@@ 3.1 Monte Carlo population synthesis code
To construct the theoretical WDLFs for the Galactic disk we made use of the Monte Carlo population synthesis code [32] for synthetic white dwarf populations. This code has been widely used in the analysis of single [e.g. 80, 81, 46, 82] and binary [e.g. 9, 18, 10, 76] populations of white dwarfs, as in studies of open and globular clusters [e.g 31, 79] and the Galactic bulge [78]. A detailed description of the code can be found in these references. Therefore, we will only review its basic inputs.
@@ 4 Impact of the axion emission in the WDLF
Figure 6: WDLFs in the $9<M_{G}<11.5$ range for different values of $g_{ae}$. In black and red dots the observed values, using Garcia-Zamora et al. DA sample [33] and the Jimenez-Esteban et al. [45] DA sample, respectively. In order to compare both samples in the same plot the counts in [45] where factorized by $N_{GZ}/N_{JE}$, with $N_{GZ}$ and $N_{JE}$ being the total number of WDs in the $9<M_{G}<11.5$ range for the [33] and [45] samples, respectively.
@@ 5 Conclusions and future work
In fact, our results are nearly a factor of two less constraining than the recent determination of [30], who analyzed white-dwarf cooling in the globular cluster 47 Tuc. This difference is consistent with our expectations: the dominant uncertainty in the Galactic-disk WDLF arises from the star-formation history, it seems natural that a stellar population with well-constrained distance, star-formation rate, and metallicity would yields stronger limits. Nevertheless, 47 Tuc represents a single cluster, and the resulting inference may be sensitive to cluster-specific systematics. In particular, the observed white-dwarf sequence is derived from specific regions of the outer field of 47 Tuc [47] and may be affected by position-dependent effects such as crowding and mass segregation. More generally, the unavoidable presence of unknown unknowns is a key consideration when assessing the robustness of astrophysical constraints on particle physics. These issues highlight the complementarity between field—and cluster—based approaches. Likewise, our bounds are comparable to, but complementary with, those derived from the TRGB [84, 11], which are subject to a very distinct set of systematic uncertainties. Fig. 8 presents these constraints in the $g_{ae}$$m_{a}$ plane, along with other bounds obtained from experimental and astrophysical observations.
Figure 8: Constraints in the $g_{ae}$$m_{a}$ plane from experimental and astrophysical observations (adapted from [59]). The orange-toned bands around $\sim 10^{-13}$ indicate the bounds derived from the Gaia DR3 WDLF (this work), red giants [11], and the recently obtained limits from Tuc 47 [30].
Besides the stellar formation rate, other limitations of this work include the neglect of mergers and binary evolution, and—on a more minor note—the adopted Galactic scale height, and the omission of axion effects on stellar thermal structure from the main sequence onward. In this context, we note that [28] report that ALP couplings can affect the estimated ages of massive CO white dwarf progenitors and modify the IFMR, potentially influencing population synthesis studies. However, these effects are not relevant for our purposes, as the disk ages explored here are substantially larger than the variations introduced by these ALP-induced corrections.
@@ Appendix A Effects of stellar formation bursts on the WDLF
For instance, in Fig. 9 we show the WDLFs in the range $9<M_{G}<13$ for different values of $g_{ae}$, assuming a constant SFR, together with their corresponding $p$-values. Although the fits systematically worsen for larger values of $g_{ae}$, none of the tests yields a good fit. Normalizing the counts over this magnitude range leads to an underestimation of the WDLF at $M_{G}\lesssim 11.5$, followed by an overestimation at the fainter end.
@@
Indeed, we find that it is possible to reproduce the observed WDLF morphology by adopting a SFR with additional degrees of freedom. In particular, we obtain acceptable fits ($p$-value $\gtrsim 74$) for a model consisting of a constant SFR over 10.6 Gyr plus two Gaussian bursts (see Fig. 10). However, this model introduces several free parameters, including the mean age, width, and normalization of each burst, as well as the normalization of the constant component. This highlights a fundamental limitation of this approach: In order to use the WDLF to robustly constrain the axion–electron coupling, the SFR must be determined independently of the white dwarf sample; otherwise, the additional free parameters could be used to artificially improve the quality of the fits, biasing the results toward certain values of $g_{ae}$. Given the current uncertainties in the observational determination of the SFR in the solar neighborhood, this requirement appears difficult to satisfy.
@@
Figure 10: Left panel: WDLF adjustment using a constant SFR over 10.6 Gyr plus two Gaussian bursts, resulting in a $p$-value $\gtrsim 74\%$. Right panel: the adjusted SFR.
@@ Acknowledgments
PIP-2971 and PICT 2020-03316. MEC acknowledges grant RYC2021-032721-I, funded by MCIN/AEI/10.13039/501100011033. MEC  and  ST  were  supported  by  the   European  UnionNextGeneration   EU/PRTR,   the   AGAUR/Generalitat   de   Catalunya   grant SGR-386/2021, and by the Spanish MINECO grant PID2023-148661NB-I00. AC is supported by an ERC STG grant (“AstroDarkLS”, grant No. 101117510). AC acknowledges the Weizmann Institute of Science for hospitality at different stages of this project and the support from the Benoziyo Endowment Fund for the Advancement of Science. MEC and ST acknowledge Jordi Isern for fruitful discussions.